OpenStudy (anonymous):

*NEED HELP* [CORRECT ME] Gayle starts to save at age 20 for an extended vacation around the world that she will take on her 45th birthday. She will contribute $1000 each year to the account, which earns 1.65% annual interest, compounded quarterly. What is the future value of this investment when she takes her trip? ----- 1000(1+0.0165)^4-1/(0.0165) ?

OpenStudy (anonymous):

Alright, well if she starts at age 20 and the trip will take place when she's 45, that's 25 years she will be adding $1000 to an account with 1.65% annual interest, compounded quarterly. Let's get some equations out for interest: Annually = P × (1 + r) = (annual compounding). Quarterly = P (1 + r/4)4 = (quarterly compounding) Now over the course of 25 years, without the interest, she has added $25,000. So to write the equation, just plug in the numbers given and solve.

OpenStudy (anonymous):

lol ohh

OpenStudy (anonymous):

A=1000*(1+0.0165) , Q=1000(1+0.0165/4)4 ?

OpenStudy (anonymous):

Good, those are your interest rates. Now, how many quarters are in 25 years?

OpenStudy (anonymous):

I dont really know :c

OpenStudy (anonymous):

A quarter is 1/4. 1 year has 4 quarters.

OpenStudy (anonymous):

I just plug (1/4) in the formula ?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

A=1000(1+0.0165/4)^(1/4) ? i'm sorry If I got it wrong i'm terrible at this